# MVPatch: More Vivid Patch for Adversarial Camouflaged Attacks on Object Detectors in the Physical World

### v2

@@ III Methodology
In this section, we first introduce the definition and characteristics of adversarial patch attacks and then detail the input and output information of threat models. Finally, we summarize the entire generation process of MVPatch in this section.
III-A Problem Formulation
The objective of this study is to generate a malicious image $\mathcal{I^{*}}$ with an adversarial patch from a benign image $\mathcal{I}$, to deceive object detection models. The complete process of generating the adversarial patch is illustrated in Fig.2. Typically, the malicious images with adversarial patches can be formulated as:
*$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P},$ (1)
where $\mathcal{M}$ represents the patch mask, which determines the size, shape, and location of the adversarial patch $\mathcal{P}$. The $\odot$ symbol denotes the Hadamard product, which multiplies the elements of the corresponding positions in the matrices $\mathcal{M}$, $\mathcal{I}$, and $\mathcal{P}$. We denote the object detection model as $f(x):x\in\mathbb{R}^{\mathbf{H\times W\times C}}\mapsto\mathcal{Y}\in\mathbb{R}^{6}$. $x$ is input image and $\mathcal{Y}$ is label of object. The generation of adversarial patches can be viewed as a constrained optimization problem:
$f(\mathcal{I^{*}})\neq f(\mathcal{I})\quad\mathrm{s.t.}\,\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P},$ (2)
where $\mathcal{P}$ represents the adversarial patch generated under the constraints of *$\lVert\mathcal{I^{*}}-\mathcal{I}\rVert\longrightarrow\mathcal{P}$, while satisfying $f(\mathcal{I^{*}})\neq f(\mathcal{I})$. Finding a closed-form solution for this optimization problem is not feasible due to the non-convex nature of the deep neural network model $f(\cdot)$. Therefore, to approximately solve this optimization problem, we utilize the following equation:
$\underset{\mathcal{P}}{\mathrm{argmax}}\,\mathcal{L}(f(\mathcal{I^{*}}),f(\mathcal{I})),$ (3)
where $\mathcal{L}$ represents the cross-entropy loss function. To solve the optimization problem, we can utilize the Adam optimizer, which is one of the most common and widely used optimization techniques.
III-B Threat Models
III-C Robustness of Adversarial Patch in the Physical Domain
Previous studies have primarily focused on applying adversarial patches in the digital domain. However, it is important to recognize that adversarial patches can also appear in the physical domain. To enhance the effectiveness of adversarial patches in physical attacks, we propose the utilization of the Non-Printable Score (NPS)[28] and Total-Variation (TV)[3] as loss functions. These will aid in the creation of more aggressive and robust adversarial patches specifically designed for physical scenarios. We present the NPS for improving the robustness when we generate adversarial patches in the physical domain using digital patterns and to ensure compatibility with the iPad color gamut and printer gamut. The formulation of the NPS is as follows:
$\mathcal{L_{\mathrm{NPS}}}=\sum_{c\in\mathcal{C},p\in\mathcal{P}}|p-c|,$ (4)
where $\mathcal{P}$ and $\mathcal{C}$ represent the pixel values of the adversarial patch and the colors that the iPad can display, respectively. Both $p$ and $c$ are elements of $\mathcal{P}$ and $\mathcal{C}$. To improve the physical robustness of adversarial patches, we utilize the Euclidean distance ||$\cdot$|| ⋅ | to constrain the generation of adversarial patches.
We also introduce the TV term, which enhances the smoothness of neighboring pixels. The TV formulation is defined as:
@@ III-C Robustness of Adversarial Patch in the Physical Domain
where $p$ denotes the pixel values of the image, while $a$ and $b$ represent the abscissa and ordinate values of a single pixel, respectively.
III-D Human-perception-based Method
There is a trade-off between the naturalness and attack performance of adversarial patches. For instance, if an adversarial patch exhibits higher attack performance, it tends to have lower naturalness. In other words, as the similarity to the source image decreases, the success rate of the attack increases. To obtain a more natural adversarial patch, we propose the human-perception-based method by utilizing the Minkowski distance as a loss function to constrain the naturalness of the patch, which we refer to as the Compared Specified Image Similarity (CSS) measurement. The CSS is formulated as follows:
$\mathcal{L_{\mathrm{CSS}}}=\left(\sum_{i=1}^{N}|\hat{x_{i}}-x_{i}|^{p}\right)^{\frac{1}{p}},$ (6)
where $\hat{x}$ represents the adversarial patch generated using the specified image $x$. We choose $p=2$ in this work and $i$ represents the number of pixels of $\hat{x}_{i}$ and $x_{i}$.
III-E Model-perception-based Method
To improve the transferability of adversarial patches, we propose the model-perception-based method. First, we present the use of two distinct loss functions to attack object detection models, as shown below:
$\mathcal{L}_{\mathrm{OBJ_{1}}}=\frac{1}{N}\sum_{i=1}^{N}S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta),$ (7)
$\mathcal{L}_{\mathrm{OBJ_{2}}}=\mathop{\max}_{\mathcal{I^{*}}}(\{S^{i}_{\mathrm{OBJ}}(\mathcal{I^{*}},\theta)\}^{N}_{i=1}),$ (8)
where $S^{i}_{\mathrm{OBJ}}$ denotes the object confidence score of the $i^{th}$ object detection model in the ensemble of object detection models (YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny). $\mathcal{I^{*}}$ is the image with adversarial patch and $\theta$ corresponds to the parameters of the object detection model.
We compare the attack performance of diverse object confidence score loss functions on various object detection models. It is observed that the adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ outperforms the one generated by $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ in terms of attacking the object detection model. The adversarial patches generated through diverse loss functions are then applied to eight object detection models (Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models) to assess their attack performance and similarity to specified images. The adversarial patch generated by $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ not only exhibits excellent attack capability (the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ is 32.20% and the average mAP of $\mathcal{L}_{\mathrm{{OBJ}_{2}}}$ is 51.80%) but also maintains a high level of naturalness. Consequently, we select $\mathcal{L}_{\mathrm{{OBJ}_{1}}}$ as our objective loss function. The results of the experiment are illustrated in Table I, where the Faster-RCNN, SSD, YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, YOLOv4-tiny, and YOLOv5 models are included as threat models. The mean mAP represents the average value of the mean Average Precision (mAP) obtained from diverse models. CS corresponds to the Cosine Similarity and ED denotes the Euclidean Distance.
III-F Dual-perception-based Framework to Generate MVPatch
To make the optimization process converge faster, we introduce additional optimization factors to the previously mentioned loss functions and formulate an ensemble loss function as Equation (9):
@@ III-F Dual-perception-based Framework to Generate MVPatch
Since optimizing the ensemble loss function with the Adam optimizer is challenging, we modified our optimization strategy. Instead of reducing the learning rate as training epochs increase, we set thresholds for the parameters and decrease the learning rate when a parameter no longer exhibits optimization during the training process. The updating strategy for the learning rate is governed by Equation (10):
$\mathcal{LR_{\mathrm{N}}}=\gamma\mathcal{LR_{\mathrm{O}}},$ (10)
@@
5:  *$\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P}$

### v3

@@ III Methodology
In this section, we provide a comprehensive overview of our Dual-Perception-Based Framework (DPBF) for generating More Vivid Patch (MVPatch) applicable to both digital and physical domains. We start by outlining the motivation behind the development of MVPatch. Following this, we introduce the foundational concepts, including notations, definitions, and problem formulation. We then detail the specifics of our proposed adversarial patch attack framework. Finally, we summarize the entire MVPatch generation process.
III-A Motivation
Adversarial Patches (APs) have gained significant attention in adversarial machine learning for their ability to manipulate object detection models. These patches can be strategically placed to cause misclassifications, posing challenges for robust detection systems. However, APs face three major challenges: Transferability, Stealthiness, and Practicality.
  • Transferability: There is insufficient theoretical support explaining why ensemble methods improve transferability, indicating a need for further research.
  • Stealthiness: GAN-based and diffusion-based methods generate realistic APs but at high computational costs and longer processing times, presenting a trade-off between realism and efficiency.
  • Practicality: In the physical domain, AP performance degrades with varying distances and angles, necessitating solutions for effective real-world deployment.
To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) designed to generate effective and stealthy adversarial patches, termed MVPatch. This framework improves the transferability of adversarial patches across various models while also enhancing their visual inconspicuousness and robustness in physical environments. We frame the adversarial transferability problem as an optimal generalization issue and introduce two modules to minimize generalization error: the Model-Perception-Based Module (MPBM), which enhances transferability through ensemble strategy, and the Human-Perception-Based Module (HPBM), which improves patch stealthiness using the Frobenius norm in Hilbert space transformations. Additionally, we incorporate regularization terms such as Total Variation (TV) and Non-Printable Score (NPS) into the DPBF to improve practical deployment in real-world scenarios. The overall framework is illustrated in Figure 2.
III-B Preliminary
Notations. We consider neural networks $\mathcal{F}(\mathcal{I},\mathcal{Y})$ for object detection tasks and let $\mathcal{I}$ be the input space of the model, and let $\mathcal{Y}=\{\mathrm{a},\mathrm{b},\mathrm{w},\mathrm{h},\mathrm{obj},\mathrm{cls}\}$. Here, $\mathrm{a}$ and $\mathrm{b}$ represent the horizontal and vertical coordinates of the detection frame, respectively. $\mathrm{w}$ and $\mathrm{h}$ denote the width and height of the detection frame. $\mathrm{obj}$ and $\mathrm{cls}$ are the object confidence scores and classification scores. We model the neural network as a mapping function $\mathcal{F}:\mathcal{I}\subseteq\mathbb{R}^{\mathbf{H\times W\times C}}\rightarrow\mathcal{Y}\subseteq\mathbb{R}^{N}$. The notations $f\in\mathcal{F}$, $x\in\mathcal{I}$, and $y\in\mathcal{Y}$ are the realizations of random variables.
For a given input $x\in\mathcal{I}$, the object detection model $f\in\mathcal{F}$ first predicts the confidence score for each label $y_{\mathrm{obj}}\in\mathcal{Y}$, denoted as $f_{\mathrm{obj}}(x)$. These confidence scores sum up to 1, i.e., $\sum_{y_{\mathrm{obj}}\in\mathcal{Y}}f_{\mathrm{obj}}(x)=1,\forall x\in\mathcal{I}$.
For a given model $f\in\mathcal{F}$, there is usually a model-dependent loss function $\mathcal{L}_{f\in\mathcal{F}}:x\in\mathcal{I}\times y\in\mathcal{Y}\rightarrow\mathbb{R}^{+}$, which is the composition of a differentiable training loss, (e.g., cross-entropy loss) $\mathcal{L}$ and model’s confidence score $f_{\mathrm{obj}}(\cdot)$. Thus, $\mathcal{L}(x,y):=\mathcal{L}(f(x),y)$, $(x,y)\in(\mathcal{I},\mathcal{Y})$.
Definition 1 (Adversarial Patch)
An adversarial patch is a perturbation $\mathcal{P}$ applied to a benign image $\mathcal{I}$ to produce a malicious image $\mathcal{I}^{*}$, with the intent of deceiving object detection models $f\in\mathcal{F}$. Formally, it can be expressed as:
$f(\mathcal{I}^{)}\neq f(\mathcal{I})=\mathcal{Y},\quad\text{where}\quad\mathcal{I}^{*}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P},$ (3)
where $\mathcal{M}$ denotes the patch mask that determines the size, shape, and placement of the adversarial patch $\mathcal{P}$. The symbol $\odot$ represents the Hadamard product, which involves element-wise multiplication of the corresponding entries in the matrices $\mathcal{M}$, $\mathcal{I}$, and $\mathcal{P}$.
Definition 2 (Generalization Error)
Consider a training dataset $\mathcal{T}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$ drawn from a joint probability distribution $P(\mathcal{I},\mathcal{Y})$. Let the hypothesis space consist of a set of models =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{M}\}$, where $M$ denotes the number of models. The generalization error $\mathcal{G}(f)$ for a model $f\in\mathcal{F}$ is defined as:
$\mathcal{G}(f)\triangleq\underbrace{\mathbb{E}\left[\mathbbm{1}\{f(x)\neq y\}\right]}_{\text{Expected Error}}-\underbrace{\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}\{f(x_{i})\neq y_{i}\}}_{\text{Empirical Error}},$ (4)
where $\mathbbm{1}\{\cdot\}$ is the indicator function that evaluates to 1 if the condition is true and 0 otherwise.
Problem Formulation. To generate adversarial patches that are both transferable and stealthy, we introduce the generalization error $\mathcal{G}(\cdot)$ and formulate the problem as an optimization task. The objective is to find an image patch $p\in\mathcal{P}$ of dimensions $(H,W,C)$, along with transformations $t\in T$, that minimize the generalization error across multiple neural networks $f\in\mathcal{F}$:
$\mathcal{P}=\operatorname*{arg\,min}_{\mathcal{P}\in\mathbb{R}^{H\times W\times C}}\mathcal{G}(f,x,p,t),\quad\text{s.t.}\quad d(t(p),t(s)),$ (5)
where $d(\cdot)$ is the distance function that quantifies the similarity between the adversarial patch $p\in\mathcal{P}$ and a specified image $s\in\mathcal{S}$. Minimizing $\mathcal{G}(f,x,p,t)$ ensures that the generated adversarial patch maintains high transferability, while satisfying $d(t(p),t(s))$ guarantees its high stealthiness.
III-C Model-Perception-Based Module
To enhance the transferability of Adversarial Patches (APs), we propose the Model-Perception-Based Module (MPBM). The MPBM incorporates generalization error analysis to derive the generalization error bounds for our method. By comparing these bounds with those of single-model and ensemble-model strategies, we divide the ensemble approach into two distinct objective functions to validate the efficacy of our method. This enables us to substantiate our approach through both theoretical derivation and empirical verification. We begin with theoretical analyses focused on generalization and stability, identifying key factors that contribute to improved transferability across various models.
Definition 3 (Ensemble Model)
Given a set of models $\{f_{i}\}^{M}_{i=1}$, ensemble strategies can be divided into averaging outputs of neural networks and maximizing outputs of neural networks. By ensembling these models, we obtain a new predictive model:
$f_{\text{ens}}(x)=\begin{cases}\frac{1}{M}\sum_{i=1}^{M}f_{i}(x)&f_{\text{ens}}^{\text{ave}}\,\,\,\text{for averaging outputs},\\ \max\{f_{i}(x)\}^{M}_{i=1}&f_{\text{ens}}^{\text{max}}\,\,\,\text{for maximizing outputs}.\end{cases}$ (6)
Remark 1
Here, $f_{\text{ens}}^{\text{max}}$ is a special case where the ensemble model strategy can also be considered a separate model. Therefore, we analyze it as an independent model in a broad sense. All subsequent references to $f_{\text{ens}}$ refer specifically to $f_{\text{ens}}^{\text{ave}}$.
III-C1 Generalization Analysis
Theorem 1 (Generalization Error of Ensemble Model)
Given an ensemble model $f_{\text{ens}}$ and a single model $f$ from the model set $\mathcal{F}$, where $(x,y)$ are drawn from the distribution $\mathcal{D}$, with $x$ and $y$ representing the input and output spaces of the models, respectively. Then,
$\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f_{\text{ens}}(x),y)\right]\leq\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f(x),y)\right].$ (7)
Proof:
$\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}(f_{\text{ens}}(x),y)\right]=\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[\mathcal{L}\left(\frac{1}{M}\sum_{i=1}^{M}f_{i}(x),y\right)\right]$ (8)
[…]
$=\mathbb{E}_{(x,y)\sim\mathcal{D};f\sim\mathcal{F}}\left[\mathcal{L}(f(x),y)\right].$ (10)
Remark 2
Since the loss function $\mathcal{L}$ is typically convex, we can utilize Jensen’s inequality to establish the relationship between the ensemble model $f_{\text{ens}}$ and single model $f$ as shown in Theorem 1. Specifically, Jensen’s inequality ensures that the loss of the ensemble model, which is a convex combination of the individual models, will not exceed the average loss of the individual models. This highlights that ensemble model generally exhibit better generalization error compared to individual models.
III-C2 Stability Analysis
Theorem 2 (Stability of the Ensemble Model)
Let Var denote variance and Cov denote covariance. Suppose the individual neural networks $f_{i},f_{j}\in\mathcal{F}$ each have a variance of $\sigma^{2}$. The variance of the ensemble model $f_{\text{ens}}$ can be expressed as:
$\text{Var}(f_{\text{ens}})=\frac{\sigma^{2}}{M}\left(1+2\rho\right),$ (11)
where $\rho=\frac{\text{Cov}(f_{i}(x),f_{j}(x))}{\sqrt{\text{Var}(f_{i}(x))\text{Var}(f_{j}(x))}}$ is the correlation coefficient between pairs of models, and $M$ is the number of models.
Proof:
$\text{Var}(f_{\text{ens}})$ $=\text{Var}\left[\frac{1}{M}\sum_{i=1}^{M}f_{i}(x)\right]$ (12)
$=\frac{1}{M^{2}}\text{Var}\left[\sum_{i=1}^{M}f_{i}(x)\right]$ (13)
$=\frac{1}{M^{2}}\underbrace{\left[\sum_{i=1}^{M}\text{Var}\left(f_{i}(x)\right)+2\sum_{i\neq j}\text{Cov}\left(f_{i}(x),f_{j}(x)\right)\right]}_{\text{Bienaymé's identity}}$ (14)
$=\frac{1}{M^{2}}\left(M\sigma^{2}+2M\rho\sigma^{2}\right)$ (15)
$=\frac{\sigma^{2}}{M}\left(1+2\rho\right).$ (16)
Remark 3
The variance of the ensemble model $f_{\text{ens}}$ is proportional to the correlation coefficient $rho$ and inversely proportional to the number of models $M$. Consequently, Eq. 11 suggests that as the number of models $M$ increases, the variance will decrease, provided that the correlation between the models $\rho$ remains low. This reduction in variance implies an enhancement in both the stability and the generalization capability of the ensemble model. Thus, the adversarial patch generation process $\hat{x}\in\mathcal{I^{*}}$ using the ensemble model $f_{\text{ens}}$ can achieve greater stability compared to a single model approach, including strategy that aims to maximize outputs in an ensemble model.
Therefore, we adopt the ensemble model strategy as outlined in Definition 3. The corresponding objective function is formulated as:
$\mathcal{L}_{\text{OBJ}}=\mathbb{E}_{f\sim\mathcal{F}}\mathbb{E}_{\hat{x}\sim\mathcal{I^{*}}}f_{\text{ens}}^{\text{ave}}(\hat{x}).$ (17)
The MPBM leverages the averaged outputs of the ensemble model $f_{\text{ens}}^{\text{ave}}$, ensuring enhanced stability and generalization during the adversarial patch generation process. By incorporating multiple models with low correlation, we achieve a more robust and reliable performance, leading to the conclusion that the ensemble strategy effectively mitigates variance and improves overall model efficacy.
III-D Human-Perception-Based Module
To enhance the stealthiness of Adversarial Patches (APs), we present the Human-Perception-Based Module (HPBM). This approach employs Kullback-Leibler (KL) divergence as a distance function and introduces a novel metric, the Compared Specified Image Similarity (CSS) measurement, to assess the similarity between the generated patch and a reference image. This methodology is analogous to the regularization techniques used in AdvART [39] and DAP [58]. While we initially considered ensemble methods, their higher computational costs compared to single models led us to adopt a more streamlined approach. Consequently, we chose a lightweight method that preserves the realism and stealthiness of APs while avoiding the computational overhead associated with ensemble strategy.
Definition 4 (Visual Reality of Adversarial Patches)
Let $\mathcal{P}$ denote the set of image patches and $\mathcal{S}$ represent the set of specified reference images. To ensure visual realism, we aim to minimize the Kullback-Leibler (KL) divergence between $\mathcal{P}$ and $\mathcal{S}$. By employing Pinsker’s inequality, the lower bound of the KL divergence can be expressed as:
$\mathcal{P}=\arg\min\mathbb{E}_{\hat{x}\sim\mathcal{P}}\left[\mathbb{KL}(\mathcal{P}||\mathcal{S})\right]\geq\frac{1}{2}\mathbb{E}_{\hat{x}\sim\mathcal{P}}\|\mathcal{P}-\mathcal{S}\|^{2}.$ (18)
Theorem 3 (Upper Bound of Generalization Error)
Consider a lower threshold of generalization error $t$ where $t\rightarrow 0$ and $t>0$. As defined in Definition 2, let $\mathcal{G}(f)$ denote the generalization error of an ensemble of $M$ models evaluated on $N$ data points. The generalization error $\mathcal{G}(f)$ can be bounded with a confidence level of $1-\gamma$, where $\gamma\in(0,1)$, as follows [35, 74]:
$\mathcal{G}(f)\leq t(M,N,\gamma)=\sqrt{\frac{1}{2N}(\log M+\log\frac{1}{\gamma})}$ (19)
Under Theorem 3, we augment the dataset with additional transformation operations to enhance data generalization. We employ the Frobenius norm in Hilbert space to design stealthy adversarial attacks, introducing the Compared Specified Image Similarity (CSS) measurement. The CSS metric is defined as follows:
$\mathcal{L}_{\text{CSS}}=\mathbb{E}_{p\sim\mathcal{P}}\mathbb{E}_{t\sim\mathcal{T}}\|t(p)-t(s)\|_{\mathcal{F}}^{2},$ (20)
where $t\in\mathcal{T}$ represents transformation operations such as rotation, flipping, and cropping.
Intuition: There exists an inherent trade-off between the naturalness and the attack efficacy of adversarial patches. Higher attack performance generally results in lower naturalness, manifesting as a reduced resemblance to the source image. Conversely, increased naturalness often corresponds to diminished attack success rates. Further details are provided in Section V-B.
III-E Practicality of Adversarial Patches in the Physical Domain
Previous research has predominantly concentrated on the application of Adversarial Patches (APs) in the digital domain. However, it is essential to recognize that APs can also be effectively utilized in physical settings. To enhance the practicality of APs for physical environments, we propose incorporating Total Variation (TV) [3] and Non-Printable Score (NPS) [29] as regularization terms. These terms facilitate the creation of more aggressive and robust adversarial patches, specifically tailored to perform effectively in real-world physical environments.
The Total Variation (TV) term is introduced to promote the smoothness of neighboring pixels. The TV regularization is defined as follows:
@@ III-E Practicality of Adversarial Patches in the Physical Domain
where $p\in\mathcal{P}$ denotes the adversarial patch, and the indices $a$ and $b$ represent the horizontal and vertical coordinates, respectively.
Additionally, we incorporate the Non-Printable Score (NPS) to enhance the robustness of adversarial patches when transitioning from digital to physical domains, ensuring compatibility with both the iPad color gamut and printer gamut. The NPS is formulated as follows:
$\mathcal{L_{\mathrm{NPS}}}=\sum_{a,b}|p_{a,b}-c_{a,b}|,$ (22)
where $p\in\mathcal{P}$ denotes the adversarial patch and $c\in\mathcal{C}$ represents the color matrix corresponding to the display capabilities of the iPad. The indices $a$ and $b$ refer to the horizontal and vertical coordinates, respectively. To further enhance the physical robustness of the adversarial patches, we employ the Euclidean distance ||$\cdot$|| ⋅ | as a constraint in the generation process.
III-F Dual-Perception-Based Framework
The Dual-Perception-Based Framework (DPBF) integrates the Model-Perception-Based Module (MPBM) and the Human-Perception-Based Module (HPBM) through the incorporation of Total Variation (TV), Non-Printability Score (NPS), and other hyperparameters, addressing the issues delineated in the Motivation section. The MPBM enhances the transferability of adversarial patches, while the HPBM ensures these patches are both realistic and stealthy. Furthermore, the inclusion of NPS and TV terms significantly augments the practicality of adversarial patches (APs), facilitating their seamless application from the digital domain to physical environments.
To expedite the optimization process, additional optimization factors have been introduced into the existing loss functions, culminating in the formulation of an aggregate loss function as presented in Eq. (23):
@@ III-F Dual-Perception-Based Framework
5:  $\mathcal{I^{*}}=(1-\mathcal{M})\odot\mathcal{I}+\mathcal{M}\odot\mathcal{P}$
@@
Since optimizing the ensemble loss function with the Adam optimizer is challenging, we modified our optimization strategy. Instead of reducing the learning rate as training epochs increase, we set thresholds for the parameters and decrease the learning rate when a parameter no longer exhibits optimization during the training process. The updating strategy for the learning rate is governed by Eq. (24):
$\mathcal{LR_{\mathrm{N}}}=\gamma\mathcal{LR_{\mathrm{O}}},$ (24)